f(x)= 1+2sinx + 3(cosx)^2, 0<=x<=2Π/3 is


a) min at x= Π/2


b)  min at x= Π/6

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Shefali Sharma,

Ans:- f(x)=1+2sinx +3cos²x

Differentiating WRT x we get = 2cosx-6sinx cosx and

the 2nd order differential is= -2sinx-6cos2x

putting dy/dx=0 we get sinx=1/3 && cosx=0

Hence applying the conditions for max and minima we get the values as

Max value:-13/3

Min value:-3

 

 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Shefali Sharma !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur